OpenStudy (anonymous):

Let f(x) =sqrt[3] x. The equation of the tangent line to f(x) at x = 8 is y =? Using this, we find our approximation for sqrt[3] {8.2} is?

OpenStudy (anonymous):

just like the last one

OpenStudy (anonymous):

i was following the same format but i keep going wrong somewhere

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x}\] right ?

OpenStudy (anonymous):

i know you take the derivative of ∛(x) which is 1/ 3∛(x)^2 then plug 8 into x to find the slope

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

and i calculated the slope would be 1/36 so the equation i got was y-0=(1/36)(x-8)

OpenStudy (anonymous):

no don't think so

OpenStudy (anonymous):

\[f'(x)=\frac{1}{3\sqrt[3]{x^2}}\]

OpenStudy (anonymous):

\[f'(8)=\frac{1}{3\sqrt[3]{8^2}}=\frac{1}{3\times 2^2}=\frac{1}{12}\]

OpenStudy (anonymous):

not sure where the 36 came from

OpenStudy (anonymous):

because i thought the whole bottom was squared but it was actually only the x

OpenStudy (anonymous):

ah right

OpenStudy (anonymous):

the equation should be y-0=1/12(x-8), correct?